OpenStudy (anonymous):

solve. 10/ 3x + 4/3 = 7 + x/2x

OpenStudy (anonymous):

This is algebra and what you want to do is solve for X. To do this, you need to move all your variables - all the numbers with the X - to one side and the integers - all the numbers without the X - to the other side. Let's start with the original equation: 3x + 7 = 2x + 10 Let's move the Xs to the left side. That way we'll end up with something that looks like X = *insert number here*. Observe that 2x is positive. To move it to the left side of the equation, you thus need to minus it from both sides. So: 3x - 2x + 7 = 2x - 2x + 10 Simplify this and the equation looks like: x + 7 = 0 + 10 The 0 on the right side? That's where your 2x was. The x on the left side? That's what happened when you moved the 2x from the right side to the left side by minusing. So now that you moved your Xs to the left side, let's move all your integers to the right side. Observe that the + 7 is a positive. To move it to the right, you need to minus it from both sides. So: x + 7 - 7 = 10 - 7 Simplified, this equation looks like: x = 3 And now, you have solved for X.

OpenStudy (ajprincess):

 Is this the question?